I need some help. I am so confused on this one. I have read my book over and ovr and it is just not clicking on how to do this.
Find all positive values for k for which each of the following can be factored.
x^2+x-k

Answer by stanbon(75887)

Find all positive values for k for which each of the following can be factored.
x^2+x-k
--------------
Assuming you mean "factoring in the Real Number System" b^2-4ac must be
greater than or equal to zero".
--------
Your problem:
a = 1; b=1; c=k
EQUATION:
1^2-4*1*k>=0
1-4k>=0
-4k>=-1
x <- 1/4
-----------------
Conclusion: x^2+x-k is factorable in the Real Number System as long
as k<=1/4
================
Cheers,
Stan H>